# Preference Elicitation for Offline Reinforcement Learning

### v1

@@ A.4.2 Analysis of the transition term
$\mathbb{E}_{d^{{\pi}}_{T}}[\tilde{R}(\tau)]-\mathbb{E}_{d^{{\pi}}_{\tilde{T}}}[\tilde{R}(\tau)]\leq HR_{max}C_{T}(\mathcal{F}_{T},\pi)\sqrt{\frac{c_{T}\log(H\mathcal{N}_{\mathcal{F}_{T}}(1/N_{o})/\delta)}{N_{o}}},$
@@
$V_{T,\tilde{R}}^{\pi}-V_{\tilde{T},\tilde{R}}^{\pi}$ $\leq R_{max}\mathbb{E}_{\tau\sim d^{\pi}_{{T}}}\left[\sum_{s_{j},a_{j}\in\tau}\|T(\cdot|s_{j},a_{j})-\tilde{T}(\cdot|s_{j},a_{j})\|_{1}\right]$
$\leq HR_{max}\mathbb{E}_{(s,a)\sim d^{\pi}_{{T}}}\left[\|T(\cdot|s,a)-\tilde{T}(\cdot|s,a)\|_{1}\right]$
$\leq HR_{max}C_{T}(\mathcal{F}_{T},\pi)\sqrt{\mathbb{E}_{(s,a)\sim D_{offline}}[\|T(\cdot|s,a)-\tilde{T}(\cdot|s,a)\|_{1}^{2}]}$
@@
$\mathbb{E}_{(s,a)\sim D_{offline}}[\|T(\cdot|s,a)-\tilde{T}(\cdot|s,a)\|_{1}^{2}]\leq\frac{1}{N_{o}}c_{T}\log(H\mathcal{N}_{\mathcal{F}_{T}}(1/N_{o})/\delta)$
@@ A.4.3 Analysis of the reward term
As in Zhan et al. [2023a], we consider the following value function: $V^{\pi}_{T,R}=\mathbb{E}_{\tau\sim d^{\pi}_{T}}[R(\tau)]-\mathbb{E}_{\tau\sim d_{pref}}[R(\tau)]$, where $d_{pref}$ is a fixed reference trajectory distribution. This baseline subtraction, which doesn’t affect either the optimal policy or the analysis of the transition term, is needed as the approximated confidence set is based on the uncertainty in preference between two trajectories, not in the reward of a single one.
@@
The concentrability coefficient w.r.t. reward classes $\mathcal{F}_{R}$, a target policy $\pi^{*}$ and a reference trajectory distribution $d_{pref}$ is defined as:
$C_{R}(\mathcal{F}_{R},\pi^{*})=\frac{\mathbb{E}_{\tau_{1}\sim d^{\pi^{*}}_{T},\tau_{2}\sim d_{pref}}\left[u_{P_{R}}(\tau_{1},\tau_{2})\right]}{\mathbb{E}_{\tau_{1},\tau_{2}\sim\mathcal{D}_{offline}}\left[u_{P_{R}}(\tau_{1},\tau_{2})\right]}$
Note that, for the purpose of our analysis, our definition differs from that of Zhan et al. [2023a] who instead consider the max ratio of difference in rewards term: $|R(\tau_{1})-R(\tau_{2})-\tilde{R}(\tau_{1})+\tilde{R}(\tau_{2})|$ over the entire function class $\mathcal{F_{R}}$.
@@
$V^{\pi^{*}}_{T,R}-V^{\pi^{*}}_{T,\hat{R}_{inf}}\leq 2\kappa C_{R}(\mathcal{F}_{T},\pi)\sqrt{\frac{c_{R}\log(\mathcal{N}_{\mathcal{F}_{R}}(1/N_{p})/\delta)}{N_{p}}},$
where $c_{R}>0$ is a constant and $\kappa=\sup_{r\in[-R_{max},R_{max}]}\frac{1}{\sigma^{\prime}(r)}$ measures the degree of non-linearity of the sigmoid function.
@@
$V^{\pi^{*}}_{T,R}-V^{\pi^{*}}_{T,\hat{R}_{inf}}$ $=\mathbb{E}_{\tau\sim d^{{\pi}^{*}}_{T}}[R(\tau)]-\mathbb{E}_{\tau\sim d_{pref}}[{R}(\tau)]-\mathbb{E}_{\tau\sim d^{{\pi}^{*}}_{T}}[\hat{R}_{inf}(\tau)]+\mathbb{E}_{\tau\sim d_{pref}}[\hat{R}_{inf}(\tau)]$
$=\mathbb{E}_{\tau_{1}\sim d^{{\pi}^{*}}_{T},\tau_{2}\sim d_{pref}}[R(\tau_{1})-R(\tau_{2})]-(\hat{R}_{inf}(\tau_{1})-\hat{R}_{inf}(\tau_{2}))]$
$\leq\kappa\mathbb{E}_{\tau_{1}\sim d^{{\pi}^{*}}_{T},\tau_{2}\sim d_{pref}}[|P_{R}(\tau_{1}\succ\tau_{2})-P_{\hat{R}_{inf}}(\tau_{1}\succ\tau_{2})|]$
$\leq\kappa\mathbb{E}_{\tau_{1}\sim d^{{\pi}^{*}}_{T},\tau_{2}\sim d_{pref}}[u_{P_{R}}(\tau_{1},\tau_{2})]$
$=\kappa C_{R}(\mathcal{F}_{T},\pi^{*})\mathbb{E}_{\tau_{1},\tau_{2}\sim\mathcal{D}_{offline}}[u_{P_{R}}(\tau_{1},\tau_{2})]$ (7)
where $\kappa=\sup_{r\in[-R_{max},R_{max}]}\frac{1}{\sigma^{\prime}(r)}$ measures the degree of non-linearity of the sigmoid function. In the first inequality, we have applied the mean value theorem, under Assumption 3.2. In the second inequality, we have used the definition of uncertainty function $u_{P_{R}}$ as we know $\hat{R}_{inf}\in\mathcal{R}$.
@@
$\mathbb{E}_{(\tau_{1},\tau_{2})\sim\mathcal{D}_{pref}}[\|P_{R}(\tau_{1}\succ\tau_{2})-P_{\tilde{R}}(\tau_{1}\succ\tau_{2})\|_{1}^{2}]\leq\frac{c_{R}\log(\mathcal{N}_{\mathcal{F}_{R}}(1/N_{p})/\delta)}{N_{p}},$ (8)
@@
$\mathbb{E}_{(\tau_{1},\tau_{2})\sim\mathcal{D}_{pref}}[u_{P_{R}}(\tau_{1},\tau_{2})]\leq 2\sqrt{\frac{c_{R}\log(\mathcal{N}_{\mathcal{F}_{g}}(1/N_{p})/\delta)}{N_{p}}}$ (9)
Under uniform sampling, the distribution of preferences in $\mathcal{D}_{pref}$ is that of the offline dataset:
$\mathbb{E}_{(\tau_{1},\tau_{2})\sim\mathcal{D}_{offline}}[u_{P_{R}}(\tau_{1},\tau_{2})]=\mathbb{E}_{(\tau_{1},\tau_{2})\sim\mathcal{D}_{pref}}[u_{P_{R}}(\tau_{1},\tau_{2})]$
@@
$V^{\pi^{*}}_{T,R}-V^{\pi^{*}}_{T,\hat{R}_{inf}}\leq 2\kappa C_{R}(\mathcal{F}_{T},\pi)\sqrt{\frac{c_{R}\log(\mathcal{N}_{\mathcal{F}_{R}}(1/N_{p})/\delta)}{N_{p}}}.$
@@
$V^{\pi^{*}}_{T,R}-V^{\pi^{*}}_{T,\hat{R}_{inf}}\leq 2\alpha\kappa C_{R}(\mathcal{F}_{T},\pi)\sqrt{\frac{c_{R}\log(\mathcal{N}_{\mathcal{F}_{R}}(1/N_{p})/\delta)}{N_{p}}},$
@@
$C_{R}^{\prime}(\mathcal{F}_{R},\pi^{*})=\frac{\mathbb{E}_{\tau_{1}\sim d^{\pi^{*}}_{T},\tau_{2}\sim d_{pref}}\left[u_{P_{R}}(\tau_{1},\tau_{2})\right]}{\mathbb{E}_{\tau_{1},\tau_{2}\sim\mathcal{D}_{pref}}\left[u_{P_{R}}(\tau_{1},\tau_{2})\right]}$
@@
$V^{\pi^{*}}_{T,R}-V^{\pi^{*}}_{T,\hat{R}_{inf}}$ $\leq\kappa\mathbb{E}_{\tau_{1}\sim d^{{\pi}^{*}}_{T},\tau_{2}\sim d_{pref}}[u_{P_{R}}(\tau_{1},\tau_{2})]$
$=\kappa C_{R}^{\prime}(\mathcal{F}_{T},\pi^{*})\mathbb{E}_{\tau_{1},\tau_{2}\sim\mathcal{D}_{pref}}[u_{P_{R}}(\tau_{1},\tau_{2})]$
$\leq 2\kappa C_{R}^{\prime}(\mathcal{F}_{T},\pi^{*})\sqrt{\frac{c_{R}\log(\mathcal{N}_{\mathcal{F}_{g}}(1/N_{p})/\delta)}{N_{p}}}$
@@
Now consider the dataset of uncertainty-sampled preferences $\mathcal{D}_{pref}$. By definition, we have:
$\mathbb{E}_{\tau_{1},\tau_{2}\sim\mathcal{D}_{pref}}\left[u_{P_{R}}(\tau_{1},\tau_{2})\right]\geq\mathbb{E}_{\tau_{1},\tau_{2}\sim\mathcal{D}_{offline}}\left[u_{P_{R}}(\tau_{1},\tau_{2})\right]$
Thus, we have: $C_{R}^{\prime}(\mathcal{F}_{T},\pi^{*})\leq C_{R}(\mathcal{F}_{T},\pi^{*})$. In other words, we can write: $C_{R}^{\prime}(\mathcal{F}_{T},\pi^{*})=\alpha C_{R}(\mathcal{F}_{T},\pi^{*})$, where $\alpha\leq 1$. This concludes our proof. ∎
@@
$V^{{\pi}^{*}}_{T,\hat{R}_{inf}}-V^{\pi^{*}}_{\hat{T}_{inf},\hat{R}_{inf}}\leq HR_{max}C_{T}(\mathcal{F}_{T},\pi^{*})\sqrt{\frac{c_{T}\log(H\mathcal{N}_{\mathcal{F}_{T}}(1/N_{o})/\delta)}{N_{o}}}$
@@
$V^{\pi^{*}}_{T,R}-V^{\pi^{*}}_{T,\hat{R}_{inf}}\leq 2\alpha\kappa C_{R}(\mathcal{F}_{T},\pi^{*})\sqrt{\frac{c_{R}\log(\mathcal{N}_{\mathcal{F}_{R}}(1/N_{p})/\delta)}{N_{p}}},$
@@ A.5.1 Suboptimality Decomposition
We decompose the suboptimality slightly differently to Equation 5, introducing the optimal offline policy (optimal in the pessimistic model under the true reward function): $\pi^{*}_{offline}=\text{argmax}_{\pi\in\Pi}V_{\hat{T}_{inf},R}^{\pi}$.
@@
$=(V^{{\pi}^{*}}_{T,R}-V^{\pi^{*}_{offline}}_{\hat{T}_{inf},\hat{R}_{inf}})-(V^{\hat{\pi}^{*}}_{T,R}-V^{{\pi}^{*}_{offline}}_{\hat{T}_{inf},\hat{R}_{inf}})$
$\leq(V^{{\pi}^{*}}_{T,R}-V^{\pi^{*}_{offline}}_{\hat{T}_{inf},\hat{R}_{inf}})-(V^{\hat{\pi}^{*}}_{T,R}-V^{\hat{\pi}^{*}}_{\hat{T}_{inf},\hat{R}_{inf}})$
$\leq V^{{\pi}^{*}}_{T,R}-V^{\pi^{*}_{offline}}_{\hat{T}_{inf},\hat{R}_{inf}}$
$=(V^{{\pi}^{*}}_{T,R}-V^{\pi^{*}}_{\hat{T}_{inf},R})+(V^{\pi^{*}}_{\hat{T}_{inf},R}-V^{\pi^{*}_{offline}}_{\hat{T}_{inf},\hat{R}_{inf}})$
$\leq\underbrace{(V^{{\pi}^{*}}_{T,R}-V^{\pi^{*}}_{\hat{T}_{inf},R})}_{\text{transition term}}+\underbrace{(V^{\pi^{*}_{offline}}_{\hat{T}_{inf},R}-V^{\pi^{*}_{offline}}_{\hat{T}_{inf},\hat{R}_{inf}})}_{\text{reward term}}$ (10)
where we have followed the same analysis as in Section A.4.1 and used the optimality of $\pi^{*}_{offline}$ in the last inequality.

### v2

@@ A.4.2 Analysis of the transition term
$\mathbb{E}_{d^{{\pi}}_{T}}[\tilde{R}(\tau)]-\mathbb{E}_{d^{{\pi}}_{\tilde{T}}}[\tilde{R}(\tau)]\leq HR_{\textrm{max}}C_{T}(\mathcal{F}_{T},\pi)\sqrt{\frac{c_{T}}{N_{o}}\log\left(\frac{H}{\delta}\mathcal{N}_{\mathcal{F}_{T}}\left(\frac{1}{N_{o}}\right)\right)},$
@@
$V_{T,\tilde{R}}^{\pi}-V_{\tilde{T},\tilde{R}}^{\pi}$ $\leq R_{\textrm{max}}\mathbb{E}_{\tau\sim d^{\pi}_{{T}}}\left[\sum_{s_{j},a_{j}\in\tau}\|T(\cdot|s_{j},a_{j})-\tilde{T}(\cdot|s_{j},a_{j})\|_{1}\right]$
$\leq HR_{\textrm{max}}\mathbb{E}_{(s,a)\sim d^{\pi}_{{T}}}\left[\|T(\cdot|s,a)-\tilde{T}(\cdot|s,a)\|_{1}\right]$
$\leq HR_{\textrm{max}}C_{T}(\mathcal{F}_{T},\pi)\sqrt{\mathbb{E}_{(s,a)\sim D_{\textrm{offline}}}[\|T(\cdot|s,a)-\tilde{T}(\cdot|s,a)\|_{1}^{2}]}$
@@
$\mathbb{E}_{(s,a)\sim D_{\textrm{offline}}}[\|T(\cdot|s,a)-\tilde{T}(\cdot|s,a)\|_{1}^{2}]\leq\frac{1}{N_{o}}c_{T}\log\left(\frac{H}{\delta}\mathcal{N}_{\mathcal{F}_{T}}\left(\frac{1}{N_{o}}\right)\right)$
@@ A.4.3 Analysis of the reward term
As in Zhan et al. [2023a], we consider the following value function: $V^{\pi}_{T,R}=\mathbb{E}_{\tau\sim d^{\pi}_{T}}[R(\tau)]-\mathbb{E}_{\tau\sim d_{\textrm{pref}}}[R(\tau)]$, where $d_{\textrm{pref}}$ is a fixed reference trajectory distribution. This baseline subtraction, which doesn’t affect either the optimal policy or the analysis of the transition term, is needed as the approximated confidence set is based on the uncertainty in preference between two trajectories, not in the reward of a single one.
@@
The concentrability coefficient w.r.t. reward classes $\mathcal{F}_{R}$, a target policy $\pi^{*}$ and a reference trajectory distribution $d_{\textrm{pref}}$ is defined as:
$C_{R}(\mathcal{F}_{R},\pi^{*})=\frac{\mathbb{E}_{\tau_{1}\sim d^{\pi^{*}}_{T},\tau_{2}\sim d_{\textup{pref}}}\left[u_{P_{R}}(\tau_{1},\tau_{2})\right]}{\mathbb{E}_{\tau_{1},\tau_{2}\sim\mathcal{D}_{\textup{offline}}}\left[u_{P_{R}}(\tau_{1},\tau_{2})\right]}$
Note that, for the purpose of our analysis, our definition differs from that of Zhan et al. [2023a] who instead consider the max ratio of difference in rewards term: $|R(\tau_{1})-R(\tau_{2})-\tilde{R}(\tau_{1})+\tilde{R}(\tau_{2})|_{1}$ over the entire function class $\mathcal{F_{R}}$.
@@
$V^{\pi^{*}}_{T,R}-V^{\pi^{*}}_{T,\hat{R}_{\textrm{inf}}}\leq 2\alpha\kappa C_{R}(\mathcal{F}_{R},\pi)\sqrt{\frac{c_{R}}{N_{p}}\log\left(\frac{1}{\delta}\mathcal{N}_{\mathcal{F}_{R}}\left(\frac{1}{N_{p}}\right)\right)},$
where $c_{R}>0$ is a constant and $\kappa=\sup_{r\in[-R_{\textrm{max}},R_{\textrm{max}}]}\frac{1}{\sigma^{\prime}(r)}$ measures the degree of non-linearity of the sigmoid function.
@@
$V^{\pi^{*}}_{T,R}-V^{\pi^{*}}_{T,\hat{R}_{\textrm{inf}}}$ $=\mathbb{E}_{\tau\sim d^{{\pi}^{*}}_{T}}[R(\tau)]-\mathbb{E}_{\tau\sim d_{\textrm{pref}}}[{R}(\tau)]-\mathbb{E}_{\tau\sim d^{{\pi}^{*}}_{T}}[\hat{R}_{\textrm{inf}}(\tau)]+\mathbb{E}_{\tau\sim d_{\textrm{pref}}}[\hat{R}_{\textrm{inf}}(\tau)]$
$=\mathbb{E}_{\tau_{1}\sim d^{{\pi}^{*}}_{T},\tau_{2}\sim d_{\textrm{pref}}}[R(\tau_{1})-R(\tau_{2})]-(\hat{R}_{\textrm{inf}}(\tau_{1})-\hat{R}_{\textrm{inf}}(\tau_{2}))]$
$\leq\kappa\mathbb{E}_{\tau_{1}\sim d^{{\pi}^{*}}_{T},\tau_{2}\sim d_{\textrm{pref}}}[|P_{R}(\tau_{1}\succ\tau_{2})-P_{\hat{R}_{\textrm{inf}}}(\tau_{1}\succ\tau_{2})|_{1}]$
$\leq\kappa\mathbb{E}_{\tau_{1}\sim d^{{\pi}^{*}}_{T},\tau_{2}\sim d_{\textrm{pref}}}[u_{P_{R}}(\tau_{1},\tau_{2})]$
$=\kappa C_{R}(\mathcal{F}_{R},\pi^{*})\mathbb{E}_{\tau_{1},\tau_{2}\sim\mathcal{D}_{\textrm{offline}}}[u_{P_{R}}(\tau_{1},\tau_{2})]$ (9)
where $\kappa=\sup_{r\in[-R_{\textrm{max}},R_{\textrm{max}}]}\frac{1}{\sigma^{\prime}(r)}$ measures the degree of non-linearity of the sigmoid function. In the first inequality, we have applied the mean value theorem, under Assumption 3.2. In the second inequality, we have used the definition of uncertainty function $u_{P_{R}}$ as we know $\hat{R}_{\textrm{inf}}\in\mathcal{R}$.
@@
$\mathbb{E}_{(\tau_{1},\tau_{2})\sim\mathcal{D}_{\textrm{pref}}}[\|P_{R}(\tau_{1}\succ\tau_{2})-P_{\tilde{R}}(\tau_{1}\succ\tau_{2})\|_{1}^{2}]\leq\sqrt{\frac{c_{R}}{N_{p}}\log\left(\frac{1}{\delta}\mathcal{N}_{\mathcal{F}_{R}}\left(\frac{1}{N_{p}}\right)\right)},$ (10)
@@
$\mathbb{E}_{(\tau_{1},\tau_{2})\sim\mathcal{D}_{\textrm{pref}}}[u_{P_{R}}(\tau_{1},\tau_{2})]\leq 2\sqrt{\frac{c_{R}}{N_{p}}\log\left(\frac{1}{\delta}\mathcal{N}_{\mathcal{F}_{R}}\left(\frac{1}{N_{p}}\right)\right)}$ (11)
Under uniform sampling, the distribution of preferences in $\mathcal{D}_{\textrm{pref}}$ is that of the offline dataset:
$\mathbb{E}_{(\tau_{1},\tau_{2})\sim\mathcal{D}_{\textrm{offline}}}[u_{P_{R}}(\tau_{1},\tau_{2})]=\mathbb{E}_{(\tau_{1},\tau_{2})\sim\mathcal{D}_{\textrm{pref}}}[u_{P_{R}}(\tau_{1},\tau_{2})]$
@@
$V^{\pi^{*}}_{T,R}-V^{\pi^{*}}_{T,\hat{R}_{\textrm{inf}}}\leq 2\kappa C_{R}(\mathcal{F}_{T},\pi)\sqrt{\frac{c_{R}}{N_{p}}\log\left(\frac{1}{\delta}\mathcal{N}_{\mathcal{F}_{R}}\left(\frac{1}{N_{p}}\right)\right)}.$
@@
$V^{\pi^{*}}_{T,R}-V^{\pi^{*}}_{T,\hat{R}_{\textrm{inf}}}\leq 2\alpha\kappa C_{R}(\mathcal{F}_{T},\pi)\sqrt{\frac{c_{R}}{N_{p}}\log\left(\frac{1}{\delta}\mathcal{N}_{\mathcal{F}_{R}}\left(\frac{1}{N_{p}}\right)\right)},$
@@
$C_{R}^{\prime}(\mathcal{F}_{T},\pi^{*})=\frac{\mathbb{E}_{\tau_{1}\sim d^{\pi^{*}}_{T},\tau_{2}\sim d_{\textrm{pref}}}\left[u_{P_{R}}(\tau_{1},\tau_{2})\right]}{\mathbb{E}_{\tau_{1},\tau_{2}\sim\mathcal{D}_{\textrm{pref}}}\left[u_{P_{R}}(\tau_{1},\tau_{2})\right]}$
@@
$V^{\pi^{*}}_{T,R}-V^{\pi^{*}}_{T,\hat{R}_{\textrm{inf}}}$ $\leq\kappa\mathbb{E}_{\tau_{1}\sim d^{{\pi}^{*}}_{T},\tau_{2}\sim d_{\textrm{pref}}}[u_{P_{R}}(\tau_{1},\tau_{2})]$
$=\kappa C_{R}^{\prime}(\mathcal{F}_{R},\pi^{*})\mathbb{E}_{\tau_{1},\tau_{2}\sim\mathcal{D}_{\textrm{pref}}}[u_{P_{R}}(\tau_{1},\tau_{2})]$
$\leq 2\kappa C_{R}^{\prime}(\mathcal{F}_{R},\pi^{*})\sqrt{\frac{c_{R}}{N_{p}}\log\left(\frac{1}{\delta}\mathcal{N}_{\mathcal{F}_{R}}\left(\frac{1}{N_{p}}\right)\right)}$
@@
Now consider the dataset of uncertainty-sampled preferences $\mathcal{D}_{\textrm{pref}}$. By definition, we have:
$\mathbb{E}_{\tau_{1},\tau_{2}\sim\mathcal{D}_{\textrm{pref}}}\left[u_{P_{R}}(\tau_{1},\tau_{2})\right]\geq\mathbb{E}_{\tau_{1},\tau_{2}\sim\mathcal{D}_{\textrm{offline}}}\left[u_{P_{R}}(\tau_{1},\tau_{2})\right]$
Thus, we have: $C_{R}^{\prime}(\mathcal{F}_{R},\pi^{*})\leq C_{R}(\mathcal{F}_{R},\pi^{*})$. In other words, we can write: $C_{R}^{\prime}(\mathcal{F}_{R},\pi^{*})=\alpha C_{R}(\mathcal{F}_{R},\pi^{*})$, where $\alpha\leq 1$. This concludes our proof. ∎
@@
$V^{{\pi}^{*}}_{T,\hat{R}_{\textrm{inf}}}-V^{\pi^{*}}_{\hat{T}_{\textrm{inf}},\hat{R}_{\textrm{inf}}}\leq HR_{\textrm{max}}C_{T}(\mathcal{F}_{T},\pi^{*})\sqrt{\frac{c_{T}}{N_{o}}\log\left(\frac{H}{\delta}\mathcal{N}_{\mathcal{F}_{T}}\left(\frac{1}{N_{o}}\right)\right)}$
@@
$V^{\pi^{*}}_{T,R}-V^{\pi^{*}}_{T,\hat{R}_{\textrm{inf}}}\leq 2\alpha\kappa C_{R}(\mathcal{F}_{R},\pi^{*})\sqrt{\frac{c_{R}}{N_{p}}\log\left(\frac{1}{\delta}\mathcal{N}_{\mathcal{F}_{R}}\left(\frac{1}{N_{p}}\right)\right)},$
@@ A.5.1 Suboptimality Decomposition
We decompose the suboptimality slightly differently to Equation 7, introducing the optimal offline policy (optimal in the pessimistic model under the true reward function): $\pi^{*}_{\textrm{offline}}=\text{argmax}_{\pi\in\Pi}V_{\hat{T}_{\textrm{inf}},R}^{\pi}$.
@@
$=(V^{{\pi}^{*}}_{T,R}-V^{\pi^{*}_{\textrm{offline}}}_{\hat{T}_{\textrm{inf}},\hat{R}_{\textrm{inf}}})-(V^{\hat{\pi}^{*}}_{T,R}-V^{{\pi}^{*}_{\textrm{offline}}}_{\hat{T}_{\textrm{inf}},\hat{R}_{\textrm{inf}}})$
$\leq(V^{{\pi}^{*}}_{T,R}-V^{\pi^{*}_{\textrm{offline}}}_{\hat{T}_{\textrm{inf}},\hat{R}_{\textrm{inf}}})-(V^{\hat{\pi}^{*}}_{T,R}-V^{\hat{\pi}^{*}}_{\hat{T}_{\textrm{inf}},\hat{R}_{\textrm{inf}}})$
$\leq V^{{\pi}^{*}}_{T,R}-V^{\pi^{*}_{\textrm{offline}}}_{\hat{T}_{\textrm{inf}},\hat{R}_{\textrm{inf}}}$
$=(V^{{\pi}^{*}}_{T,R}-V^{\pi^{*}}_{\hat{T}_{\textrm{inf}},R})+(V^{\pi^{*}}_{\hat{T}_{\textrm{inf}},R}-V^{\pi^{*}_{\textrm{offline}}}_{\hat{T}_{\textrm{inf}},\hat{R}_{\textrm{inf}}})$
$\leq\underbrace{(V^{{\pi}^{*}}_{T,R}-V^{\pi^{*}}_{\hat{T}_{\textrm{inf}},R})}_{\text{transition term}}+\underbrace{(V^{\pi^{*}_{\textrm{offline}}}_{\hat{T}_{\textrm{inf}},R}-V^{\pi^{*}_{\textrm{offline}}}_{\hat{T}_{\textrm{inf}},\hat{R}_{\textrm{inf}}})}_{\text{reward term}}$ (12)
where we have followed the same analysis as in Section A.4.1 and used the optimality of $\pi^{*}_{\textrm{offline}}$ in the last inequality.